# Passive harmonic mode-locked laser on lithium niobate integrated photonics

### v1

@@ Abstract
Mode-locked lasers (MLLs) are essential for a wide range of photonic applications, such as frequency metrology, biological imaging, and high-bandwidth coherent communications. The growing demand for compact and scalable photonic systems is driving the development of MLLs on various integrated photonics material platforms. Along these lines, developing MLLs on the emerging thin-film lithium niobate (TFLN) platform holds the promise to greatly broaden the application space of MLLs by harnessing TFLN ’s unique electro-optic (E-O) response and quadratic optical nonlinearity. Here, we demonstrate the first electrically pumped, self-starting passive MLL in lithium niobate integrated photonics based on its hybrid integration with a GaAs quantum-well gain medium and saturable absorber. Our demonstrated MLL generates 4.3-ps optical pulses centered around 1060 nm with on-chip peak power exceeding 44 mW. The pulse duration can be further compressed to 1.75 ps via linear dispersion compensation. Remarkably, passive mode-locking occurs exclusively at the second harmonic of the cavity free spectral range, exhibiting a high pulse repetition rate $\sim$20 GHz. We elucidate the temporal dynamics underlying this self-starting passive harmonic mode-locking behavior using a traveling-wave model. Our work offers new insights into the realization of compact, high-repetition-rate MLLs in the TFLN platform, with promising applications for monolithic ultrafast microwave waveform sampling and analog-to-digital conversion.
Figure 1: Hybrid GaAs–TFLN passive harmonic MLL a, Schematic of the passive harmonic MLL. The GaAs RSOA chip contains an SOA section and an SA section, and it is butt-coupled to a TFLN extended cavity to form a Fabry-Perot laser cavity. When passive mode-locking occurs, two ultrashort light pulses exist within one cavity round-trip. b, Optical microscope image of the TFLN nanophotonic waveguide forming a passive extended cavity. Scale bar: 100 µm. c, Butt-coupling interface between the GaAs RSOA chip and the TFLN chip. d, Fundamental TE mode in the GaAs waveguide at 1060 nm. e Fundamental TE mode in the input TFLN waveguide at 1060 nm. f, SEM image of the CDC in the broadband Sagnac loop mirror. Scale bar: 2 µm. The coupler has a total coupling length of 54 µm and a gap of 600 nm between the waveguide top surfaces. g, Simulated (blue solid line) and measured (red symbols) reflection spectrum of the Sagnac loop mirror.
@@
Despite recent advances in actively MLLs in TFLN 40, the realization of high-repetition-rate, passive MLLs in TFLN remains elusive, yet warrants investigation for several reasons. First, an intra-cavity passive mode-locking element, such as a saturable absorber, allows for the generation of shorter pulses than its active counterparts due to the significantly shorter net gain windows41, 42. Second, a passive MLL eliminates the need for external high-frequency microwave drivers, thereby significantly reducing system complexity. Passive mode-locking can also overcome the bandwidth limitation of a microwave driver, allowing for the generation of a high repetition rate pulse train. In this work, we present the first electrically pumped, self-starting passive harmonic MLL on TFLN. The MLL consistently and exclusively exhibits mode-locking at $\sim$20 GHz, which is the second harmonic of the cavity free spectral range (FSR). Importantly, such a harmonic passive mode-locking behavior enables a substantial increase in repetition rate without requiring cavity length reduction, which is highly favorable for developing high-repetition-rate on-chip MLLs. We also develop a traveling-wave model combined with a fast–slow carrier reservoir framework to elucidate the complex interplay between gain and saturable absorber dynamics, which underlies the observed self-starting passive harmonic mode-locking behavior.
@@
As shown in Fig. 4a, the temporal pulse train captured by the fast photodetector and the high-speed oscilloscope directly confirms the second harmonic passive mode locking at 20.1 GHz repetition rate. To gain deeper insight into the harmonic mode-locking dynamics, we employed a traveling wave model that separates the laser into two distinct regions: an active region comprised of SOA, SA, and the gap sections; and a passive TFLN extended cavity, which only imposes linear dispersion and linear loss. Detailed parameters of our model are shown in the Supplementary Information Section 6. The temporal and spatial evolutions of the optical fields in both SOA and SA are governed by the partial differential equations54:
$\small\frac{\partial A^{\pm}(z,t)}{\partial t}\pm\frac{\partial A^{\pm}(z,t)}{\partial z}=j\frac{\omega_{\mathrm{0}}v_{\mathrm{g}}\Gamma}{2n_{\text{eff}}c}\chi(\omega,N_{\mathrm{f}})A^{\pm}(z,t)-\beta\,A^{\pm}(z,t),$ (1)

### v2

@@ Abstract
Mode-locked lasers (MLLs) are essential for a wide range of photonic applications, such as frequency metrology, biological imaging, and high-bandwidth coherent communications. The growing demand for compact and scalable photonic systems is driving the development of MLLs on various integrated photonics material platforms. Along these lines, developing MLLs on the emerging thin-film lithium niobate (TFLN) platform holds the promise to greatly broaden the application space of MLLs by harnessing TFLN ’s unique electro-optic (E-O) response and quadratic optical nonlinearity. Here, we demonstrate the first electrically pumped, self-starting passively MLL in lithium niobate integrated photonics based on its hybrid integration with a GaAs quantum-well gain medium and saturable absorber. Our demonstrated MLL generates 4.3-ps optical pulses centered around 1060 nm with on-chip peak power exceeding 44 mW. The pulse duration can be further compressed to 1.75 ps via linear dispersion compensation. Remarkably, passive mode-locking occurs exclusively at the second harmonic of the cavity free spectral range, exhibiting a high pulse repetition rate $\sim$20 GHz. We elucidate the temporal dynamics underlying this self-starting passive harmonic mode-locking behavior using a traveling-wave model. Our work offers new insights into the realization of compact, high-repetition-rate MLLs in the TFLN platform, with promising applications for monolithic ultrafast microwave waveform sampling and analog-to-digital conversion.
Figure 1: Hybrid GaAs–TFLN passive harmonic MLL a, Schematic of the passive harmonic MLL. The GaAs RSOA chip contains an SOA section and an SA section, and it is butt-coupled to a TFLN extended cavity to form a Fabry-Perot laser cavity. When passively mode-locking occurs, two ultrashort light pulses exist within one cavity round-trip. b, Optical microscope image of the TFLN nanophotonic waveguide forming a passive extended cavity. Scale bar: 100 µm. c, Butt-coupling interface between the GaAs RSOA chip and the TFLN chip. d, Fundamental TE mode in the GaAs waveguide at 1060 nm. e Fundamental TE mode in the input TFLN waveguide at 1060 nm. f, SEM image of the CDC in the broadband Sagnac loop mirror. Scale bar: 2 µm. The coupler has a total coupling length of 54 µm and a gap of 600 nm between the waveguide top surfaces. g, Simulated (blue solid line) and measured (red symbols) reflection spectrum of the Sagnac loop mirror.
@@
Despite recent advances in actively MLLs in TFLN 40, the realization of high-repetition-rate, passively MLLs in TFLN remains elusive, yet warrants investigation for several reasons. First, an intra-cavity passive mode-locking element, such as a saturable absorber, allows for the generation of shorter pulses than its active counterparts due to the significantly shorter net gain windows41, 42. Second, a passively MLL eliminates the need for external high-frequency microwave drivers, thereby significantly reducing system complexity. Passive mode-locking can also overcome the bandwidth limitation of a microwave driver, allowing for the generation of a high repetition rate pulse train. In this work, we present the first electrically pumped, self-starting passive harmonic MLL on TFLN. The MLL consistently and exclusively exhibits mode-locking at $\sim$20 GHz, which is the second harmonic of the cavity free spectral range (FSR). Importantly, such a harmonic passive mode-locking behavior enables a substantial increase in repetition rate without requiring cavity length reduction, which is highly favorable for developing high-repetition-rate on-chip MLLs. We also develop a traveling-wave model combined with a fast–slow carrier reservoir framework to elucidate the complex interplay between gain and saturable absorber dynamics, which underlies the observed self-starting passive harmonic mode-locking behavior.
@@
As shown in Fig. 4a, the temporal pulse train captured by the fast photodetector and the high-speed oscilloscope directly confirms the second harmonic passive mode-locking at 20.1 GHz repetition rate. To gain deeper insight into the harmonic mode-locking dynamics, we employed a traveling wave model that separates the laser into two distinct regions: an active region comprised of SOA, SA, and the gap sections; and a passive TFLN extended cavity, which only imposes linear dispersion and linear loss. Detailed parameters of our model are shown in the Supplementary Information Section 6. The temporal and spatial evolutions of the optical fields in both SOA and SA are governed by the partial differential equations54:
$\frac{\partial A^{\pm}(z,t)}{\partial t}\pm\frac{\partial A^{\pm}(z,t)}{\partial z}=j\frac{\omega_{\mathrm{0}}v_{\mathrm{g}}\Gamma}{2n_{\text{eff}}c}\chi(\omega,N_{\mathrm{f}})A^{\pm}(z,t)-\beta\,A^{\pm}(z,t),$ (1)